# Predicting solvation free energies with an implicit solvent machine learning potential

### v1

@@ Results
The parity plot (Fig. 2a) also highlights the FreeSolv’s uneven distribution with respect to the hydration free energy value. There are only a few molecules with large and negative hydration free energy. The largest two (Mobley ID 9534740 and D-mannitol) are particularly problematic for the Amber and CHARMM models. These molecules, discussed further in the next section, are part of a general trend. The MAE increases with decreasing hydration free energy (Fig. 2d) for all models. However, ReSolv model is the most robust model in this respect. See Supplementary Figure 2 for point-wise results and Supplementary Figure 1 for the training dataset results. For Amber and CHARMM models, a possible explanation could be an inaccurate partial charge assignment. As shown by Jämbeck et al. [66], the hydration free energy is very sensitive to the choice of charge computation method and can differ by several kcal/mol depending on the method used. We found that the absolute sum of the Gasteiger partial charge per atom is highly correlated with the hydration free energy as is the polar surface area of the molecules (Supplementary Figure 3). Indeed, keeping in mind the net-neutrality of the molecules, the absolute Gasteiger partial charge per atom is sensitive to local charge gradients contributed by polar fragments within the molecule, whence its correlation with polar surface area. The computed Pearson correlation coefficient is r=-0.71, which is comparable to the correlation with the related polar surface area (r=-0.74) but significantly larger than correlation with other molecular properties such as dipole moment (r=-0.42) or volume of a solute (r=-0.18). In particular, we find that the absolute sum of the charges increases with decreasing hydration free energy. An accurate partial charge assignment is, therefore, particularly important for molecules with large negative hydration free energies, potentially explaining why the predictions for classical force fields tend to worsen with decreasing hydration free energy. With this insight, we depict in Supplementary Figure 4 the difference between the $U_{\mathrm{sol}}$ and $U_{\mathrm{vac}}$ atomic contributions to the potential energy for a fixed configuration and find a correspondence with the Gasteiger partial charges from the heavy atoms. The energy differences highlight that the changes are predominant around polar atoms, agreeing with what we would expect based on the above correlations, supporting that the learning of $U_{\mathrm{sol}}$ is physically informed.
@@
Figure 3: Error analysis. The box plots of the mean absolute error (MAE) with respect to (a) chemical functional groups, (b) number of heavy atoms, and (c) number of heavy atom types are shown for the test set molecules. The red line denotes the median, red dashed line mean, the box ranges are from the first to the third quartile, and the whiskers correspond to the 1.5x interquartile range. The gray bar plots depict the number of corresponding samples in the training set. We compare three different models: ReSolv (blue), Amber (GAFF; red), and CHARMM (CGenFF; green). In subplot a, the ‘other’ group corresponds to the remaining functional groups with less than three samples in the test set.
@@
First, we examine the error with respect to chemical functional groups. The chemical space covered by the FreeSolv database is quite extensive given the relatively small size of the database [31, 13, 14]. ReSolv demonstrates uniform error across the chemical functionalities, which is, to some extent, expected given the training set construction. However, low MAE is also found for the functional groups rarely appearing in the training set, i.e., for the ‘other’ group where we merge functional groups with sparse occurrence in the train and test sets. In contrast, Amber and CHARMM models exhibit larger fluctuations in MAE for some functional groups (e.g., primary alcohol for Amber or primary amine for CHARMM) displaying notably higher errors. Note that the same conclusion can be drawn for the molecules in the training set (Supplementary Figure 5). Next, we consider the models’ robustness with respect to the size of the molecule. From 14 heavy atoms onwards, an increased MAE can be seen for all models, most notably for the Amber model. Lastly, in terms of heavy atom types, all three predictive models display larger errors for molecules with four heavy atom types. However, we attribute this result mainly to the small sample size containing problematic molecules, i.e., the MAE is computed on only four molecules with the nitralin molecule (discussed in the next section) contributing most to the error. The absence of a similar trend for the molecules in the training set (Supplementary Figure 5) further supports our claim.
Error correlations
@@
We noticed several outliers in the parity plot (Fig. 2a) for which the predictions are largely off for all three investigated models. Error correlation between different modeling approaches can be used to identify possible erroneous data points. Previous outlier analysis stimulated corrections of some experimental data points that are part of the FreeSolv database [67]. For example, the initially provided experimental value for D-mannitol was -27.79 kcal/mol which was later corrected to -23.62 kcal/mol. Nevertheless, doubts about the validity of the experimental value remained as the COSMO-RS prediction deviated by 6 kcal/mol [67]. As a result, some recent studies excluded D-mannitol from the test set [18, 17]. Interestingly, ReSolv’s prediction of D-mannitol’s hydration free energy is -27.42, in excellent agreement with the original experimental value.
@@
We compute the Pearson correlation coefficient of signed and absolute errors to investigate the error correlation between ReSolv, Amber, and CHARMM predictions (Supplementary Figure 6). Overall, we find positive correlations between models, with the largest correlation of 0.6 between the Amber and CHARMM models. In Fig. 4, we show point-wise error correlations and mark the molecules with large and correlated errors (exact values are given in Supplementary Table 1).
@@
Figure 4: Error correlations. Correlation of errors in kcal/mol between the ReSolv, Amber (GAFF), and CHARMM (CGenFF) models for the test dataset. The gray dash-dotted line indicates the perfect correlation. Eight molecules with absolute errors greater than 1.7 kcal/mol simultaneously for x- and y-axes are marked with distinct symbols as shown in the legend. In Supplementary Table 1 we list the mean errors across all models and experimental uncertainties for these eight molecules.
@@
DISCUSSION
@@
This work presents the Solvation Free Energy Path Reweighting (ReSolv) to efficiently learn an ML-based potential energy surface in an implicit solvent by utilizing experimental solvation free energy data and reweighting in training. Since the framework was showcased with QM7-X and FreeSolv datasets, the trained ReSolv model enables an accurate hydration free energy prediction for small organic compounds, which can be directly utilized in, e.g., de novo drug design. However, the spanned chemical space of molecules could be enlarged with other datasets. In addition, the same methodology can be used to obtain an ML potential for implicit solvents other than water. For instance, extensively available octanol-water partition data could be used to derive a model in an implicit octanol environment. Predicting lipophilicity, correlated with oil-water partition coefficient, is also a critical parameter in the pharmaceutical industry as drug candidates must be sufficiently lipophilic to penetrate the lipid core of membranes but not too lipophilic that they remain there [68].
@@
For the FreeSolv database, ReSolv’s MAE is close to experimental uncertainty. The obtained accuracy is also similar to the previously reported structure-property models, employing ML algorithms to directly predict hydration free energy from the physics-inspired fingerprints [17, 8, 29, 18] or molecular structure [19, 20, 26, 25, 22, 23, 9, 24]. The latter employ graph-based neural network architectures and typically perform better than the former with the reported MAE in the range of 0.58-0.76 and RMSE in the range of 0.82-1.23. This results indicate that further improvements will likely require an enlarged and improved experimental database rather than a better modeling approach or neural network architecture. In line with this conclusion is a recent study [21] achieving a notably lower MAE of 0.42 where a graph neural network was first pre-trained on a large dataset and later fine-tuned on the FreeSolv database, i.e., by exploiting the transfer learning approach. ReSolv showcases good generalization across the functional space but a decreased generalization to larger molecules, which should be considered when planning future experiments. In addition, we found error correlations between our models and classical atomistic models that were not trained on the FreeSolv database. Molecules with large and correlated errors, such as nitralin and pirimor, call for a reevaluation of the experimental data. These points should be excluded when training future models to avoid detrimental effects.
@@
In a broader context, ReSolv could be a first step towards a general implicit solvent ML potential with better accuracy and efficiency than classical atomistic models. Solvation free energy encodes solute-solvent interactions, which in turn govern many biomolecular processes, including folding, aggregation, and ligand binding. In vivo, these processes occur in an aqueous solution rendering the hydration free energy a general benchmark for classical force field validation [69, 10, 1], fine-tuning [70, 2], comparison [71, 72], and calibration [73, 74]. Nevertheless, before using the ReSolv model in general simulations, the model needs to be tested and potentially additionally trained on other available experimental data, which we leave for future studies. Similar as for other implicit solvent ML potentials, using a prior potential (i.e., a fixed simple potential) is likely also necessary to achieve stable simulations of large macromolecules [58]. Concerning computational efficiency, the ReSolv and ML-based implicit solvent models, in general, provide a speed-up compared to the classical explicit solvent force fields due to reduced degrees of freedom and associated accelerated dynamics [75], even though evaluating an ML potential requires much more operations than an empirical potential [76]. Further gains could be obtained with constrained dynamics or a coarse-grained representation of solute molecules, enabling larger integration timesteps.
METHODS
Both vacuum and solvent ML potentials are based on the GNN architecture NequIP [63] as implemented in JAX MD [77]. The architectural hyperparameters of NequIP are listed in Supplementary Table 2.
where $\gamma_{U}$ and $\gamma_{F}$ are weighting factors, $N_{bs}$ is the batch size, $N_{i,a}$ the number of atoms of molecule $i$, and $k$ iterates over the x-, y-, z-dimensions. $U_{i}$ is the energy of the $i$-th molecule in the batch, and $F_{ijk}$ is the force in direction k of atom j from the i-th molecule in the batch. Subscript DFT denotes the ab initio data target. We normalize the target energies and forces. The energies are shifted by the mean energy of the training data and scaled with the average root mean squared force of the training data. The forces are scaled with the average root mean square force of the training data. Further, we do not employ learnable scaling or shifting and set the per-atom scaling to 1 and the per-atom shift to 0. To pick the best model, we employ early stopping on the validation dataset. For the numerical optimization hyperparameters, see Supplementary Table 3.
In the second stage, we employ the RDKit [78] to generate the 3D structure of a molecule based on the SMILES string provided by the FreeSolv database. Next, we perform the energy minimization of structures with the MMFM94 force field [79, 80, 81, 82, 83]. The obtained configurations are used to run a 300 ps initial equilibration simulation followed by a 200 ps production run for each molecule. All simulations are performed in the NVT ensemble and numerically integrated with the velocity Verlet scheme using a 1 fs timestep. To match the conditions of the experimental data the temperature is kept at 298.15 K with the Langevin thermostat with damping factor set to 1/ps. We check the stability of our simulations by evaluating that the last configuration of the molecules can be represented by a single graph and that no atoms diverged from the molecule. In training, new simulations are initialized with the last configuration of the previous simulation and consist of 50 ps equilibration and a 200 ps production run. We sample every 5 ps during the production run, i.e., 40 samples per trajectory. The reweighting effective sample size threshold is fixed to $\bar{N}_{\mathrm{eff}}=0.9$. The numerical optimization hyperparameters are reported in Supplementary Table 3.
The reported hydration free energies are computed with the BAR method using only the end states, i.e., vacuum and water states. As previously reported, intermediate states are not necessary for implicit solvent models due to the sufficient overlap between the vacuum and solvated ensembles of the solute [33]. We first run 300 ps of equilibration and sample every 5 ps in the subsequent 200 ps production run. We perform convergence tests (Supplementary Figure 7) for 12 randomly selected molecules in the test set to ensure sufficient sampling.
Computations were performed with double precision on Nvidia A100 80GB GPUs.
Data availability
The data will be made publicly available at https://github.com/tummfm/ReSolv upon acceptance of the paper.
The QM7-X dataset by Hoja et al. [54] is available at https://doi.org/10.5281/zenodo.4288677.
The FreeSolv dataset by Mobley and Guthrie [13] is available at https://escholarship.org/uc/item/6sd403pz.
Code availability
The code will be made publicly available at https://github.com/tummfm/ReSolv.git upon acceptance of the paper.
@@ METHODS
J.Z. conceptualized the study. S.R. and A.F.B. implemented and applied the ReSolv method and conducted simulations and postprocessing. All authors planned the study, analyzed and interpreted the results, and wrote the paper.
Acknowledgements
@@
Competing interests
The authors declare no competing interests.

### v2

@@ III.1 ReSolv outperforms classical explicit solvent force fields on FreeSolv database
The parity plot (Fig. 2a) also highlights the FreeSolv’s uneven distribution with respect to the hydration free energy value. There are only a few molecules with large and negative hydration free energy. The largest two (Mobley ID 9534740 and D-mannitol) are particularly problematic for the Amber and CHARMM models. These molecules, discussed further in the next section, are part of a general trend. The MAE increases with decreasing hydration free energy (Fig. 2d) for all models. However, ReSolv model is the most robust model in this respect. See supplementary material Fig. S3 for point-wise results and supplementary material Fig. S1 for the training dataset results. For Amber and CHARMM models, a possible explanation could be an inaccurate partial charge assignment. As shown by Jämbeck et al. Jämbeck et al. (2013), the hydration free energy is very sensitive to the choice of charge computation method and can differ by several kcal/mol depending on the method used. We found that the absolute sum of the Gasteiger partial charge per atom is highly correlated with the hydration free energy as is the polar surface area of the molecules (supplementary material Fig. S4). Indeed, keeping in mind the net-neutrality of the molecules, the absolute Gasteiger partial charge per atom is sensitive to local charge gradients contributed by polar fragments within the molecule, whence its correlation with polar surface area. The computed Pearson correlation coefficient is r=-0.71, which is comparable to the correlation with the related polar surface area (r=-0.74) but significantly larger than correlation with other molecular properties such as dipole moment (r=-0.42) or volume of a solute (r=-0.18). In particular, we find that the absolute sum of the charges increases with decreasing hydration free energy. An accurate partial charge assignment is, therefore, particularly important for molecules with large negative hydration free energies, potentially explaining why the predictions for classical force fields tend to worsen with decreasing hydration free energy. With this insight, we depict in supplementary material Fig. S5 the difference between the $U_{\mathrm{sol}}$ and $U_{\mathrm{vac}}$ atomic contributions to the potential energy for a fixed configuration and find a correspondence with the Gasteiger partial charges from the heavy atoms. The energy differences highlight that the changes are predominant around polar atoms, agreeing with what we would expect based on the above correlations, supporting that the learning of $U_{\mathrm{sol}}$ is physically informed.
@@
Figure 3: Error analysis. The box plots of the mean absolute error (MAE) with respect to (a) chemical functional groups, (b) number of heavy atoms, and (c) number of heavy atom types are shown for the test set molecules. The red line denotes the median, red dashed line mean, the box ranges are from the first to the third quartile, and the whiskers correspond to the 1.5x interquartile range. The gray bar plots depict the number of corresponding samples in the training set. We compare three different models: ReSolv (blue), Amber (GAFF; red), and CHARMM (CGenFF; green). In subplot (a), the ‘other’ group corresponds to the remaining functional groups with less than three samples in the test set.
@@
First, we examine the error with respect to chemical functional groups. The chemical space covered by the FreeSolv database is quite extensive given the relatively small size of the database Karwounopoulos et al. (2023); Mobley and Guthrie (2014); Duarte Ramos Matos et al. (2017). ReSolv demonstrates uniform error across the chemical functionalities, which is, to some extent, expected given the training set construction. However, low MAE is also found for the functional groups rarely appearing in the training set, i.e., for the ‘other’ group where we merge functional groups with sparse occurrence in the train and test sets. In contrast, Amber and CHARMM models exhibit larger fluctuations in MAE for some functional groups (e.g., primary alcohol for Amber or primary amine for CHARMM) displaying notably higher errors. Note that the same conclusion can be drawn for the molecules in the training set (supplementary material Fig. S6). Next, we consider the models’ robustness with respect to the size of the molecule. From 14 heavy atoms onwards, an increased MAE can be seen for all models, most notably for the Amber model. Lastly, in terms of heavy atom types, all three predictive models display larger errors for molecules with four heavy atom types. However, we attribute this result mainly to the small sample size containing problematic molecules, i.e., the MAE is computed on only four molecules with the nitralin molecule (discussed in the next section) contributing most to the error. The absence of a similar trend for the molecules in the training set (supplementary material Fig. S6) further supports our claim.
III.2 ReSolv generalizes well to unseen functional groups
The results thus far demonstrate ReSolv’s ability to accurately predict the hydration free energies of molecules with chemical functional groups seen during training. Given the typically limited availability of training samples, it is especially desirable for ML potentials to extrapolate effectively into unseen chemical space. To assess ReSolv’s chemical generalizability, we employ the same training procedure and dataset prescribed in Sect. III.1, but with all samples of a given functional group (including multifunctional molecules) removed from the training set and compare the performance to the cases when seen.
We conducted two such trainings, excluding samples with either primary amine or primary alcohol functionalities. These functional groups were selected due to their challenging nature for CHARMM and Amber, respectively (see Fig. 3a). With the test sets identical, we found that the errors associated with the unseen functional groups were highly consistent with those observed when the same groups were seen in training, see Fig. 4. This consistency demonstrates ReSolv’s robustness and capacity to predict well even for unencountered regions in chemical space.
Figure 4: Extrapolation to unseen functional groups. Box plots of the mean absolute error (MAE) with respect to chemical functional groups, comparing four different models: ReSolv with primary amines unseen during training (dark blue); ReSolv with primary alcohols unseen during training (light blue); Amber (GAFF; red); and CHARMM (CGenFF; green). The ‘other’ group corresponds to the remaining functional groups with less than three samples in the test set. The red line denotes the median, red dashed line mean, the box ranges are from the first to the third quartile, and the whiskers correspond to the 1.5x interquartile range.
III.3 Error correlations
@@
We noticed several outliers in the parity plot (Fig. 2a) for which the predictions are largely off for all three investigated models. Error correlation between different modeling approaches can be used to identify possible erroneous data points. Previous outlier analysis stimulated corrections of some experimental data points that are part of the FreeSolv database Reinisch and Klamt (2014). For example, the initially provided experimental value for D-mannitol was -27.79 kcal/mol which was later corrected to -23.62 kcal/mol. Nevertheless, doubts about the validity of the experimental value remained as the COSMO-RS prediction deviated by 6 kcal/mol Reinisch and Klamt (2014). As a result, some recent studies excluded D-mannitol from the test set Zhang et al. (2023); Riniker (2017). Interestingly, ReSolv’s prediction of D-mannitol’s hydration free energy is -27.42 kcal/mol, in excellent agreement with the original experimental value.
@@
We compute the Pearson correlation coefficient of signed and absolute errors to investigate the error correlation between ReSolv, Amber, and CHARMM predictions (supplementary material Fig. S7). Overall, we find positive correlations between models, with the largest correlation of 0.6 between the Amber and CHARMM models. In Fig. 5, we show point-wise error correlations and mark the molecules with large and correlated errors (exact values are given in supplementary material Table S1).
@@
Figure 5: Correlation of errors in kcal/mol between the ReSolv, Amber (GAFF), and CHARMM (CGenFF) models for the test dataset. The gray dash-dotted line indicates the perfect correlation. Eight molecules with absolute errors greater than 1.7 kcal/mol simultaneously for x- and y-axes are marked with distinct symbols as shown in the legend. In supplementary material Table 1 we list the mean errors across all models and experimental uncertainties for these eight molecules.
@@
III.4 ReSolv achieves four orders of magnitude speedup compared to the explicit solvent ML potentials
Lastly, we compare the ReSolv’s performance to explicit solvent ML potentials. Using the same ML potential architecture, implementation, and common simulation setup in the literature, the measured computational cost differs by four orders of magnitude (supplementary material). The ReSolv’s speed up is due to due to two factors: (i) faster execution of MD step due to reduced number of particles, (ii) reduced number of required MD steps due to enhanced sampling.
Surprisingly, ReSolv’s substantial computational gains do not come at the cost of accuracy. To the contrary, comparison with the recent work Moore, Cole, and Csanyi (2024) on hydration free energy predictions for six selected molecules in the FreeSolv database suggests that the ReSolv’s predictions are in better agreement with the experimental results (Fig. 6). More precisely, ReSolv model achieves better accuracy for five out of six molecules, including ethane, which is part of our test dataset, while the other five molecules were included in the training dataset.
Figure 6: Comparison with an explicit solvent ML potential. The hydration free energy prediction for six molecules in the FreeSolv database recently tested by Moore et al. Moore, Cole, and Csanyi (2024) with an explicit solvent ML potential. Apart from caffeine, ReSolv’s predictions are closer to experimental results. This includes ethane which is part of our test dataset.
IV Conclusion
@@
This work presents the Solvation Free Energy Path Reweighting (ReSolv) to efficiently learn an ML-based potential energy surface in an implicit solvent by utilizing experimental solvation free energy data and reweighting in training. Since the framework was showcased with QM7-X and FreeSolv datasets, the trained ReSolv model enables an accurate hydration free energy prediction for small organic compounds, which can be directly utilized in, e.g., de novo drug design. However, the spanned chemical space of molecules could be enlarged with other datasets. In addition, the same methodology can be used to obtain an ML potential for implicit solvents other than water. For instance, extensively available octanol-water partition data could be used to derive a model in an implicit octanol environment. Predicting lipophilicity, correlated with oil-water partition coefficient, is also a critical parameter in the pharmaceutical industry as drug candidates must be sufficiently lipophilic to penetrate the lipid core of membranes but not too lipophilic that they remain there Van de Waterbeemd, Lennernäs, and Artursson (2004). Transfer learning strategy would likely accelerate the learning of other solvents. It is reasonable to assume that similar solvents modify the interactions similarly. For example, the knowledge gained by learning an implicit water model could benefit the model for other polar solvents, such as ethanol. This setting could be particularly useful for solvents for which solvation free energy data is scarce.
@@
For the FreeSolv database, ReSolv’s MAE is close to experimental uncertainty. The obtained accuracy is also similar to the previously reported structure-property models, employing ML algorithms to directly predict hydration free energy from the physics-inspired fingerprints Riniker (2017); Hutchinson and Kobayashi (2019); Weinreich, Browning, and von Lilienfeld (2021); Zhang et al. (2023) or molecular structure Lim and Jung (2021); Low, Coote, and Izgorodina (2022); Chen et al. (2021a); Pathak, Mehta, and Priyakumar (2021); Wu et al. (2018); Yang et al. (2019); Lim and Jung (2019); Cho and Choi (2019). The latter employ graph-based neural network architectures and typically perform better than the former with the reported MAE in the range of 0.58-0.76 kcal/mol and RMSE in the range of 0.82-1.23 kcal/mol. This results indicate that further improvements will likely require an enlarged and improved experimental database rather than a better modeling approach or neural network architecture. In line with this conclusion is a recent study Zhang, Xia, and Zhang (2022) achieving a notably lower MAE of 0.42 kcal/mol where a graph neural network was first pre-trained on a large dataset and later fine-tuned on the FreeSolv database, i.e., by exploiting the transfer learning approach. ReSolv showcases good generalization across the functional space but a decreased generalization to larger molecules, which should be considered when planning future experiments. In addition, we found error correlations between our models and classical atomistic models that were not trained on the FreeSolv database. Molecules with large and correlated errors, such as nitralin and pirimor, call for a reevaluation of the experimental data. These points should be excluded when training future models to avoid detrimental effects.
@@
In a broader context, ReSolv could be a first step towards a general implicit solvent ML potential with better accuracy and efficiency than classical atomistic models. Solvation free energy encodes solute-solvent interactions, which in turn govern many biomolecular processes, including folding, aggregation, and ligand binding. In vivo, these processes occur in an aqueous solution rendering the hydration free energy a general benchmark for classical force field validation Duarte Ramos Matos, Calabro, and Mobley (2019); Nicholls et al. (2008); Fossat, Zeng, and Pappu (2021), fine-tuning Huang et al. (2017); Nerenberg et al. (2012), comparison Kashefolgheta et al. (2020); Shirts et al. (2003), and calibration Oostenbrink et al. (2004); Kashefolgheta and Verde (2017). Nevertheless, before using the ReSolv model in general simulations, the model needs to be tested and potentially additionally trained on other available experimental data, which we leave for future studies. Similar as for other implicit solvent ML potentials, using a prior potential (i.e., a fixed simple potential) is likely also necessary to achieve stable simulations of large macromolecules Durumeric et al. (2023). Concerning computational efficiency, the ReSolv and ML-based implicit solvent models, in general, provide a speed-up compared to the classical explicit solvent force fields due to reduced degrees of freedom and associated accelerated dynamics Meinel and Müller-Plathe (2020); Katzberger and Riniker (2024), even though evaluating an ML potential requires much more operations than an empirical potential Charron et al. (2023). Further gains could be obtained with constrained dynamics or a coarse-grained representation of solute molecules, enabling larger integration timesteps.
Supplementary material
See the supplementary material for hyperparameters and additional results including performance and error analysis on train dataset, absolute error vs experimental hydration free energy, correlation of molecular properties with experimental hydration free energy, potential energy difference vs Gasteiger charges, Pearson correlation of errors, convergence of the hydration free energy computation.
@@
J.Z. conceptualized the study. S.R. and A.F.B. contributed equally to this work. S.R. and A.F.B. implemented and applied the ReSolv method and conducted simulations and postprocessing. All authors planned the study, analyzed and interpreted the results, and wrote the paper.
Conflicts of interest
There are no conflicts to declare.
Acknowledgements.
@@
DATA AVAILABILITY STATEMENT
The data that support the findings of this study are available within the article and its supplementary material and will be made openly available at https://github.com/tummfm/ReSolv upon acceptance of the paper. The QM7-X dataset by Hoja et al. Hoja et al. (2021) is available at https://doi.org/10.5281/zenodo.4288677. The FreeSolv dataset by Mobley and Guthrie Mobley and Guthrie (2014) is available at https://escholarship.org/uc/item/6sd403pz. The code will be made publicly available at https://github.com/tummfm/ReSolv.git upon acceptance of the paper.